You have 7 sweaters that you want to take on vacation but only have enough room to pack 4 sweaters in your suitcase. How many different combinations of sweaters can you bring?
Answer by Maths68(1474)

Use
nCr=n!/r!(n-r)!
Given
n=7, r=4
nCr=n!/r!(n-r)!
nCr=7!/4!(7-4)!
nCr=7!/4!3!
nCr=7*6*5*4!/4!3!
Cancel our 4!
nCr=7*6*5/3!
nCr=7*6*5/3*2*1
nCr=210/6
nCr=35
35 Combinations.